OpenStudy (anonymous):

Simplify the expression. 2[(13-1)^2/2] (A) 12 (B) 24 (C) 72 (D) 144

JoelTheBoss (joel_the_boss):

ok first you need to simplify 13−1 to 12\[2 \times \frac{ 12^2 }{ 2}\]

JoelTheBoss (joel_the_boss):

Next Simplify 12^2 to 144\[2 \times \frac{ 144 }{ 2 }\] Cross multiply and since the denominator is the same it stays 2.\[\frac{ 288 }{ 2}\] Then you just go ahead and simplify and you get \[288\div2= 144~~~< Answer :) \]

JoelTheBoss (joel_the_boss):

Did this help?

JoelTheBoss (joel_the_boss):

@confluxepic Was I right?

OpenStudy (confluxepic):

Yes. But next time try to guide the user.

OpenStudy (confluxepic):

@JesseCole618 Do you understand how @Joel_the_boss got it.

JoelTheBoss (joel_the_boss):

I don't think he wants my help.. v_v xD And I did guide him. I showed my steps.

OpenStudy (confluxepic):

P: Parenthesis. E: Exponents. M: Multiplication. D: Division. A: Addition. S: Subtraction.

OpenStudy (anonymous):

@Joel_the_boss It doesn't matter. He/She can just ignore the steps and take the answer. Does that make sense?

JoelTheBoss (joel_the_boss):

Yes. :) Thanks guys. :P